a line is drawn on a graph using the points (-1.2) and (5.7) as shown. which of the following expressions represents the slope of the line?
Found 3 solutions by ikleyn, greenestamps, math_tutor2020:
Answer by ikleyn(53942)

.

The set of following expressions is empty in this post.


So, the correct answer is: each of the following expressions.




Answer by greenestamps(13367)



"(-1.2)" and "(5.7)" are two points on a number line -- not two points in an xy plane. Assuming the number line is horizontal, the slope of the line between the two points is 0.

Use commas instead of periods to show the coordinates of points in two dimensions: (-1,2) and (5,7)


Answer by math_tutor2020(3838)


If you meant to say the points (-1, 2) and (5, 7), then you can use the slope formula as shown below.

m = slope
m = (change in y)/(change in x)
m = (y2 - y1)/(x2 - x1)
m = (7-2)/(5 - (-1))
m = (7-2)/(5 + 1)
m = 5/6